OpenStudy (anonymous):

how would I solve a negative quadratic function? example: f(x) = -x^2+6x-5

OpenStudy (anonymous):

\[ -x^2+6x-5=0\iff x^2-6x+5=0\] solve the "positive" one

OpenStudy (anonymous):

same way you would a positive one. only A=-1 insead of 1. or you could just multiply both sides by -1 to make the leading coefficient to positive.

hero (hero):

First, set f(x) = 0: 0 = -x^2 + 6x - 5 Next factor out the negative: 0 = -(x^2 - 6x + 5) Then divide both sides by -1: 0 = x^2 - 6x + 5

hero (hero):

@Satellite73 is such a sniper.

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

(x-5)(x-1) divide both by -1 = x=-5 x=-1?

hero (hero):

After dividing both sides by -1, you don't have to do that process anymore.

OpenStudy (anonymous):

your factoring is right. just dont re divide.

hero (hero):

You only do it once, then solve the quadratic as you normally do.

OpenStudy (anonymous):

if you did redivide it would be -x=-5 though keep that in mind

hero (hero):

(x - 5)(x - 1) = 0 x - 5 = 0 x - 1 = 0 Then solve from there.

OpenStudy (anonymous):

ohh i see thank you guys :D